# Foundation Models in Medical Imaging - A Review and Outlook

### v3

@@ 1 Introduction
Medical image analysis plays a crucial role in modern healthcare, with techniques such as X-ray, MRI, and microscopic analysis of tissue samples providing essential insights for diagnosis, treatment planning, and disease monitoring. However, a significant challenge in this field is the scarcity of labeled data, which is often expensive and time-consuming to obtain due to privacy concerns, the need for expert annotation, and the complexity of acquiring high-quality labels. Traditional supervised learning methods, which rely heavily on large amounts of manually annotated data, struggle to perform effectively in such settings.
Foundation Models (FMs), a recent development within Artificial Intelligence (AI),offer a promising solution to this challenge. Unlike traditional supervised models, FMs are pre-trained on large, diverse, unlabeled datasets to learn rich, general representations of the data. These representations capture broad patterns and features, enabling FMs to achieve strong performance across a wide range of tasks. When fine-tuned on smaller, task-specific labeled datasets, FMs not only reduce the dependency on large annotated datasets but also often outperform traditional methods, thanks to the depth and generality of their pre-trained knowledge.
Despite their potential, the use of FMs in medical image analysis is still in its early stages. This review aims to provide a detailed examination of the application of foundation models in medical image analysis, focusing specifically on three primary imaging domains: pathology, radiology, and ophthalmology. We found that the wide range of imaging modalities discussed in the literature can be broadly categorized into these three application groups. As a result, our review is concentrated on these domains to provide an analysis representative of imaging techniques in the field. By reviewing over 150 studies, we offer a thorough overview of the current state of research and the various models developed and employed across these domains.
We focus on FMs whose purpose is medical image analysis, considering only those that process visual information, or a combination of vision and language, and exclude language-only models. By providing a comprehensive analysis and summarizing landmark works, we aim to equip researchers and practitioners with the knowledge needed to navigate this rapidly evolving field.
@@ 1.1 Overview
To establish a solid foundation for our review, Section 2 provides an overview of foundation models and the typical pipeline for deployment via three core concepts: large-scale architectures, self-supervised learning, and adaptation to downstream tasks. This section equips readers with the background necessary to understand the subsequent analyses. Subsequently, sections 3, 4, and 5 provide reviews of foundation models in pathology, radiology, and ophthalmology, respectively. These include introductions into the different imaging domains and concise table overviews summarizing landmark works. Additionally, Section 6 identifies and discusses key challenges and open questions in applying FMs to medical image analysis. By outlining these challenges, we aim to guide future research and innovation in the field.
@@
  • Foundation Model Background: We offer an in-depth explanation of foundation models and how their pipelines function, dividing it into three main building blocks.
  • Comprehensive Review: We examine over 150 studies on foundation models in medical image analysis and provide tables summarizing the various models used, serving as a quick reference for readers.
  • Modality-Specific Organization: We categorize our review by the main imaging domains used in clinical practice (pathology, radiology, and ophthalmology) for detailed insights.
  • Identification of Challenges: We highlight key challenges and open questions in the field, directing future research efforts.
@@ 2 Background
The term Foundation Model (FoMo/FM) was popularized by the Stanford Institute for Human-Centered Artificial Intelligence, defining them as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks” [1].
@@
The development of foundation models has been significantly influenced by the realization that larger datasets lead to improved model performance. This understanding is supported by the concept of scaling laws, which describe predictable improvements in performance as model size, dataset size, and computational resources are increased [2]. Specifically, as datasets grow larger and model parameters and computational resources are scaled accordingly, models generally exhibit enhanced performance [3]. In the general domain, FM network architectures can have trainable parameters in the range of billions, and dataset sizes are also substantial, ranging from tens of thousands to hundreds of millions of images. For example, OpenAI’s CLIP model was trained on 400 million image-text pairs [4]. Given the limitations of obtaining large labeled datasets, especially in specialized fields like medical imaging, the focus has shifted toward utilizing self-supervised learning (SSL) methods to leverage large amounts of unlabeled data, thus leading to the current foundation models paradigm.
@@
The scale of foundation models not only contributes to their generalizability but can also lead to the model exhibiting novel behaviors and insights that smaller models might not demonstrate. Such emergent abilities include improved understanding, generation, and contextualization of data, which can significantly enhance model performance across various tasks [5].
The initial success of Large Language Models (LLMs), such as the GPT series from OpenAI [6] and BERT from Google [7], marked the start of foundation models. Building on their success in natural language processing, the concept expanded into computer vision with the development of Vision Foundation Models (VFMs) and Vision-Language Foundation Models (VLFMs), which integrate both visual and textual data.
@@
FMs are particularly useful in medical image analysis because they address the limitations of traditional computer vision approaches. Traditionally, computer vision models rely on supervised learning with labeled datasets, which are scarce in medical settings due to the time, cost, and expertise required for annotation. To mitigate this, models are often pre-trained on datasets like ImageNet, which contain natural images. However, ImageNet-trained models are limited by their focus on natural images, which differ significantly from medical data in terms of texture, structure, and context. As a result, they often struggle to generalize across diverse medical specialties and imaging domains.
FMs, on the other hand, are pre-trained on vastly larger and more diverse datasets, using self-supervised learning techniques. Unlike traditional pre-training, which is often constrained to a single domain (e.g., natural images from ImageNet), FMs are designed to learn universal representations that capture a broader range of patterns and features. This makes them more adaptable to the unique characteristics of medical data, even when fine-tuned on limited labeled datasets. By leveraging their ability to generalize across domains and tasks, FMs excel at a wide range of medical imaging applications, including lesion segmentation, disease classification, generating medical summaries from imaging data and patient records, and answering clinical questions related to images, such as estimating tumor sizes or predicting disease progression.
@@ 2.1 FM Core Technical Concepts
Foundation models are characterized by three core concepts that set them apart from traditional deep learning methods and allow them to excel across various domains. At the foundation of these models’ capabilities are: 1. large-scale pretraining on extensive datasets, enabled by 2. self-supervised learning objectives that extract meaningful patterns without manual annotation. Through this process, the models learn rich, generalizable representations in the form of embeddings: high-dimensional vectors that capture clinically relevant patterns and relationships in the data. Unlike traditional deep learning models designed for specific tasks, foundation models organize knowledge in these versatile embeddings, which can then be 3. efficiently adapted to various downstream clinical applications through multiple approaches. Below, we examine these concepts in detail.
@@ 2.1.1 Large-Scale Pre-Training
The purpose of large-scale encoders in foundation models is to extract meaningful patterns from large datasets and convert raw data into generalizable embeddings. These architectures can be broadly categorized based on the data type they handle. For textual data, foundation models typically utilize transformer-based architectures, which excel at processing sequences and capturing contextual relationships within text. For handling visual data, there is a choice between using Convolutional Neural Networks (CNNs) or Vision Transformers (ViTs). These architectures will be discussed below.
@@
CNNs have been fundamental in computer vision and are known for their ability to extract hierarchical features from image data. They assume that nearby pixels are related (locality) and that different parts of an image are processed similarly (weight sharing). These inductive biases help CNNs learn effectively from limited training data, making them particularly efficient for tasks where local patterns, such as edges and textures, are crucial. ResNet [8], for example, is a popular choice for CNN-based foundation models due to its depth and residual connections, which help maintain the flow of information in deeper architectures. ConvNeXt [9] is another modern adaptation of CNNs that incorporates some elements of transformer architectures to improve performance while retaining the beneficial inductive biases of CNNs.
@@
The transformer architecture was originally developed for natural language processing [10] and later successfully adapted to visual data through vision transformers [11]. Rather than using sliding convolutional filters, as CNNs do, ViTs split an image into fixed-size patches (for example 16×16 pixels) and treat each patch as an individual element in a sequence. These patches are first converted into feature vectors through a linear projection. Crucially, since this process loses spatial information, learned positional encodings are added to preserve where each patch was located in the original image. The architecture’s core innovation lies in its self-attention mechanism, which allows every patch to dynamically interact with every other patch, determining how much attention to pay to different parts of the image based on their relevance. This global attention mechanism allows the model to capture both local patterns and long-range dependencies equally effectively.
@@
Despite their differences, CNNs and transformers can be powerful backbone architectures for foundation models in various domains. CNNs leverage inductive biases like locality and weight sharing, allowing them to effectively extract spatial features and perform well with limited data. These biases make CNNs efficient for tasks where local patterns (e.g., edges, textures) are crucial, but they can limit the model’s ability to capture long-range dependencies, which are often important in large-scale or complex datasets. In contrast, ViTs have minimal inductive biases, enabling them to model global dependencies across images. This flexibility allows ViTs to excel in large-scale pretraining, particularly with self-supervised learning, although it also requires significantly larger training datasets to achieve optimal performance [12]. As a result, ViTs are more commonly used as backbones for foundation models due to their scalability. However, CNNs remain effective in scenarios with limited data and are often integrated into hybrid architectures, such as Pyramid Vision Transformers (PvT), Swin Transformers, and Convolutional Vision Transformers (CvT) [13, 14, 15], which combine the strengths of CNNs and transformers to enhance data efficiency and multiscale feature representation.
@@ 2.1.2 Self-Supervised Learning
Discriminative SSL methods can be further divided based on their underlying techniques. One of the most prominent approaches within this category is contrastive learning, which uses the InfoNCE loss [16] to align augmented views of the same image (positive samples) and differentiate between different images (negative samples). Prominent frameworks such as SimCLR, MoCo, and SwAV [17, 18, 19] gained popularity for effectively utilizing unlabeled data with pretext tasks, like identifying matching augmented views. On the other hand, distillation-based methods like BYOL, DINO, and DINOv2 [20, 21, 22] eliminate the need for negative samples by training a student network using "soft labels" from a teacher network.
@@
While discriminative methods were most commonly used in computer vision, the introduction of vision transformer in 2021 marked a significant advancement in computer vision’s approach to self-supervision [23]. Like its natural language counterpart, masked language modeling (MLM), vision transformers used masked or altered tokenized image patches to learn to predict the missing or altered parts rather than merely distinguishing between samples. This approach, known as Masked Image Modeling (MIM), has since led to the development of various generative methods in computer vision.
One of the earlier notable examples of MIM methods include BEiT, which uses vector quantization to map continuous image patches to discrete visual tokens. It does not predict pixel values directly but learns to predict these discrete tokens, which are created from image patches [24]. Initially relying on an external, offline-trained tokenizer, BEiT’s efficiency was later improved by iBOT with the introduction of an online tokenizer [25].
Another recent but popular approach within generative SSL is Masked Autoencoder (MAE), which employs an asymmetric autoencoder framework. MAE’s unique feature is its encoder, which processes only unmasked patches of an image while ignoring masked ones. A lightweight decoder is then tasked with reconstructing the original image from these encoded representations and the masked patches. By focusing only on reconstructing the missing (masked) parts, the decoder encourages the model to learn meaningful visual representations. This design not only reduces computational overhead but also ensures the model captures semantic content efficiently, making MAE faster and more effective than previous generative frameworks [26].
More recently, the image-based implementation of the Joint Embedding Predictive Architecture (I-JEPA) [27] offers a novel SSL approach by predicting embeddings of masked regions based on context from unmasked regions within the same image. While similar in concept to MAE, I-JEPA focuses on predicting representations in the embedding space rather than in the image space. Utilizing vision transformers for encoding, I-JEPA enhances semantic understanding and computational efficiency, avoiding the pixel-level details typical of generative approaches.
@@
While most self-supervised learning objectives are designed for single-stream data, multimodal foundation models that handle both vision and language require specialized approaches to manage these diverse data types. Combining multiple modalities can be handled at various stages in the foundation model pipeline, a spectrum known as early (or feature level) to late fusion (or decision level).
Late fusion involves using separate encoders for each data type, each with its own SSL objectives. These modalities are processed independently and combined later in the pipeline, as seen, for example, in [28]. This method offers flexibility, allowing for the tailored processing of each modality and simplified integration at a later stage.
In contrast, early fusion integrates features from different modalities before or during the representation learning process. This results in a unified representation space that captures interactions between data types. Early fusion models, such as CLIP [4] and ALIGN [29], often utilize Image-Text Contrastive Learning (ITC). ITC aligns paired images and texts by maximizing their similarity and minimizing the similarity between mismatched pairs, effectively representing related multimodal data closely in a shared embedding space. Additionally, ITC can leverage unpaired data by aligning images with separate texts describing similar concepts, using techniques like self-distillation [30].
Finally, mid-fusion, and multi-layer fusion approaches, fall in between early and late fusion. These methods combine modalities at different preprocessing stages, often by creating layers within the neural network specifically designed for data fusion purposes [31].
The choice between early and late fusion methods depends on the specific requirements of the task: early fusion captures fine-grained interactions between modalities but may involve higher dimensionality. Late fusion allows for greater flexibility in unimodal encoder pipeline design but may fail to capture complex interaction patterns. Mid fusion offers a balance by integrating modalities at specific stages, though it can lead to increased model complexity.

### v4

@@ 1 Introduction
Medical image analysis plays a crucial role in modern healthcare, with techniques such as X-ray, magnetic resonance imaging, and microscopic tissue analysis providing essential insights for diagnosis, treatment planning, and disease monitoring. A persistent challenge in this field is the scarcity of labeled data, which is expensive and time-consuming to obtain due to privacy concerns, expert annotation requirements, and the complexity of acquiring high-quality labels. Traditional supervised learning methods, which rely heavily on large annotated datasets, often struggle under these constraints.
Foundation Models (FMs), a recent development within Artificial Intelligence (AI), offer a promising solution. Pre-trained on large, diverse, unlabeled datasets, FMs learn rich, general representations that capture broad patterns and features. Fine-tuning these models on smaller, task-specific labeled datasets reduces dependency on extensive annotations while often improving performance compared to traditional approaches.
Despite their potential, the application of FMs in medical image analysis is still emerging. This review focuses on vision-based FMs for medical imaging across three primary domains: pathology, radiology, and ophthalmology. By systematically analyzing over 150 studies, we provide a comprehensive overview of the current state of research and the various models developed and employed across these domains.
@@ 1.1 Overview
Section 2 introduces foundation models and their typical deployment pipeline, covering three core concepts: large-scale architectures, self-supervised learning, and adaptation to downstream tasks. Sections 3, 4, and 5 review FMs in pathology, radiology, and ophthalmology, respectively, including domain introductions and tables summarizing landmark works. Section 6 discusses key challenges and open questions to guide future research.
@@
  • Foundation Model Background: In-depth explanation of FMs and their pipeline divided into three main building blocks.
  • Comprehensive Review: Examination of over 150 studies with summary tables for quick reference.
  • Modality-Specific Organization: Review organized by the main imaging domains (pathology, radiology, ophthalmology) for detailed insights.
  • Identification of Challenges: Highlighting key challenges and open questions to guide future research.
1.2 Relation to Other Works
Several recent reviews have addressed foundation models in healthcare. Some take a broad perspective, covering applications beyond imaging, such as natural language processing for electronic health records, drug discovery, medical robotics, and education [zhang2023challenges, qiu2023large, khan2024comprehensive, zhang2024data]. Others concentrate primarily on large language models, without substantial coverage of vision-based foundation models [liu2024large, yuan2023large, zhou2023survey, he2023survey, thirunavukarasu2023large]. Within medical imaging, azad2023foundational provide a general overview but do not analyze individual studies in detail, while paschali2025foundation focus specifically on radiology and emphasize overarching concepts rather than surveying the breadth of implementations.
Our review differs in two key ways. First, we focus exclusively on vision-based foundation models for medical image analysis, excluding language-only models and broader healthcare applications. Second, we provide both a wide coverage and a structured analysis: over 150 studies are systematically reviewed, organized by imaging domain (pathology, radiology, and ophthalmology), and presented with summary tables and detailed discussion. This structure allows us to place individual contributions in context, highlight methodological trends and domain-specific differences, and identify open challenges. In this way, our work complements prior reviews by offering a comprehensive and modality-specific resource for researchers in medical imaging.
@@ 2 Background
The term Foundation Model was popularized by the Stanford Institute for Human-Centered Artificial Intelligence, defining them as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks” [bommasani_opportunities_2022].
@@
The development of foundation models has been significantly influenced by the realization that larger datasets lead to improved model performance. This understanding is supported by the concept of scaling laws, which describe predictable improvements in performance as model size, dataset size, and computational resources are increased [kaplan2020scaling]. Specifically, as datasets grow larger and model parameters and computational resources are scaled accordingly, models generally exhibit enhanced performance [achiam2023gpt]. In the general domain, FM network architectures can have trainable parameters in the range of billions, and dataset sizes are also substantial, ranging from tens of thousands to hundreds of millions of images. For example, OpenAI’s CLIP model was trained on 400 million image-text pairs [radford2021learning]. Given the limitations of obtaining large labeled datasets, especially in specialized fields like medical imaging, the focus has shifted toward utilizing self-supervised learning (SSL) methods to leverage large amounts of unlabeled data, thus leading to the current foundation models paradigm.
@@
The scale of foundation models not only contributes to their generalizability but can also lead to the model exhibiting novel behaviors and insights that smaller models might not demonstrate. Such emergent abilities include improved understanding, generation, and contextualization of data, which can significantly enhance model performance across various tasks [awais_foundational_2023].
The initial success of Large Language Models (LLMs), such as the GPT series from OpenAI [brown_language_2020] and BERT from Google [devlin_bert_2019], marked the start of foundation models. Building on their success in natural language processing, the concept expanded into computer vision with the development of Vision Foundation Models (VFMs) and Vision-Language Foundation Models (VLFMs), which integrate both visual and textual data.
@@
Traditionally, deep learning models for image analysis are trained using supervised learning, which requires large amounts of labeled data. In medical imaging, such annotated datasets are often scarce due to the time, cost, and expertise required for labeling, which severely limits the performance of deep learning models. Some methods try to overcome this limitation by pre-training on large labeled natural image datasets, such as ImageNet, which are more readily available. However, because natural images differ greatly from medical images in texture, structure, and context, these models often struggle to generalize across medical specialties and imaging modalities.
Foundation Models overcome this limitation by learning from large, diverse, unlabeled datasets using self-supervised learning. Instead of being trained for a single task, they learn general representations that capture broad patterns and features in the data. These rich representations can then be fine-tuned for a variety of medical imaging tasks—such as classification, segmentation, detection, and diagnosis—using relatively few labeled examples. This approach reduces the reliance on annotated datasets while improving generalization across imaging modalities and clinical domains.
@@ 2.1 FM Core Technical Concepts
Figure 1: Overview of core technical concepts behind foundation models for medical image analysis. Dotted lines represent optional components. The backbone architecture consists of vision encoders processing imaging data, while auxiliary encoders incorporate complementary modalities in multimodal settings, such as clinical notes, patient records, or molecular data. Self-supervised learning methods, such as contrastive, generative, and distillation approaches, are used to pretrain the model and learn rich representations from the data. In multimodal scenarios, representations from different encoders can be fused before adaptation to downstream tasks. Adaptation strategies range from lightweight, zero/few-shot methods to heavier approaches such as full fine-tuning.
Foundation models are characterized by three core concepts that set them apart from traditional deep learning methods. At the foundation of these models’ capabilities are: 1. large-scale pretraining on extensive datasets, enabled by 2. self-supervised learning objectives that extract meaningful patterns without manual annotation. Through this process, the models learn rich, generalizable representations in the form of embeddings: high-dimensional vectors that capture clinically relevant patterns and relationships in the data. Unlike traditional deep learning models designed for specific tasks, foundation models organize knowledge in these versatile embeddings, which can then be 3. efficiently adapted to various downstream clinical applications through multiple approaches (see Figure 1). Below, we examine these concepts in detail.
@@ 2.1.1 Large-Scale Pre-Training
The purpose of large-scale encoders in foundation models is to extract meaningful patterns from large datasets and convert raw data into generalizable embeddings. Depending on the data type, different architectures are used: transformer-based models are typically employed for textual data, while visual data are processed using either Convolutional Neural Networks (CNNs) or Vision Transformers (ViTs). In this review, we focus on the latter two families, which form the backbone of vision-based foundation models.
@@
CNNs have been fundamental in computer vision and are known for their ability to extract hierarchical features from image data. They assume that nearby pixels are related (locality) and that different parts of an image are processed similarly (weight sharing). These inductive biases help CNNs learn effectively from limited training data, making them particularly efficient for tasks where local patterns, such as edges and textures, are crucial. ResNet [resnet], for example, is a popular choice for CNN-based foundation models due to its depth and residual connections, which help maintain the flow of information in deeper architectures. ConvNeXt [liu2022convnet] is another modern adaptation of CNNs that incorporates some elements of transformer architectures to improve performance while retaining the beneficial inductive biases of CNNs.
@@
The transformer architecture was originally developed for natural language processing [vaswani2017attention] and later successfully adapted to visual data through ViTss [dosovitskiy2021image]. Rather than using sliding convolutional filters, as CNNs do, ViTs split an image into fixed-size patches (for example 16×16 pixels) and treat each patch as an individual element in a sequence. These patches are first converted into feature vectors through a linear projection. Crucially, since this process loses spatial information, learned positional encodings are added to preserve where each patch was located in the original image. The architecture’s core innovation lies in its self-attention mechanism, which allows every patch to dynamically interact with every other patch, determining how much attention to pay to different parts of the image based on their relevance. This global attention mechanism allows the model to capture both local patterns and long-range dependencies equally effectively.
@@
Despite their differences, CNNs and transformers can be powerful backbone architectures for foundation models in various domains. CNNs leverage inductive biases like locality and weight sharing, allowing them to effectively extract spatial features and perform well with limited data. These biases make CNNs efficient for tasks where local patterns (e.g., edges, textures) are crucial, but they can limit the model’s ability to capture long-range dependencies, which are often important in large-scale or complex datasets. In contrast, ViTs have minimal inductive biases, enabling them to model global dependencies across images. This flexibility allows ViTs to excel in large-scale pretraining, particularly with self-supervised learning, although it also requires significantly larger training datasets to achieve optimal performance [goldblum2024battle]. As a result, ViTs are more commonly used as backbones for foundation models due to their scalability. However, CNNs remain effective in scenarios with limited data and are also integrated into hybrid architectures, such as Pyramid Vision Transformers (PvT), Swin Transformers, and Convolutional Vision Transformers (CvT) [wang2021pyramid, liu2021swin, wu2021cvt], which combine the strengths of CNNs and transformers to enhance data efficiency and multiscale feature representation.
@@ 2.1.2 Self-Supervised Learning
Discriminative SSL methods can be further divided based on their underlying techniques. One of the most prominent approaches within this category is contrastive learning, which uses the InfoNCE loss [oord2018representation] to align augmented views of the same image (positive samples) and differentiate between different images (negative samples). Well-known examples of contrastive learning frameworks include SimCLR, MoCo, and SwAV [chen2020simple, he_momentum_2020, caron_unsupervised_2021]. On the other hand, distillation-based methods like BYOL, DINO, and DINOv2 [grill2020bootstrap, caron_emerging_2021, oquab_dinov2_2024] eliminate the need for negative samples by training a student network using "soft labels" from a teacher network.
@@
While computer vision has traditionally focused on discriminative learning approaches, the introduction of ViTs in 2021 marked an important shift in self-supervised representation learning [balestriero_cookbook_2023]. By representing images as patch tokens,analogous to word tokens in natural language processing, ViTs enabled the transfer of language-based pretraining strategies to vision. In natural language processing, transformers are commonly trained with Masked Language Modeling (MLM), where parts of the input text are hidden and must be predicted. This inspired Masked Image Modeling (MIM), where subsets of image patches are masked or altered and the model learns to reconstruct them rather than solely distinguishing between samples. MIM has since driven substantial advances in generative self-supervised methods for vision.
One of the earlier notable MIM approaches is BEiT, which applies vector quantization to map continuous image patches to discrete visual tokens. Unlike approaches that predict pixel values directly, BEiT predicts these discrete tokens [bao_beit_2022]. BEiT initially relied on an externally pre-trained tokenizer, but iBOT improved the method by introducing an online tokenizer, which improved efficiency[zhou_ibot_2022].
Another recent but popular approach within generative SSL is Masked Autoencoder (MAE), which employs an asymmetric autoencoder framework. MAE’s unique feature is its encoder, which processes only unmasked patches of an image while ignoring masked ones. A lightweight decoder is then tasked with reconstructing the original image from these encoded representations and the masked patches. By focusing only on reconstructing the missing (masked) parts, the decoder encourages the model to learn meaningful visual representations. This design not only reduces computational overhead but also ensures the model captures semantic content efficiently, making MAE faster and more effective than previous generative frameworks [he2021masked].
More recently, the image-based implementation of the Joint Embedding Predictive Architecture (I-JEPA) [assran2023self] offers a novel SSL approach by predicting embeddings of masked regions based on context from unmasked regions within the same image. While similar in concept to MAE, I-JEPA predicts representations in the embedding space rather than the image space. Operating in the embedding space reduces sensitivity to pixel-level details typical of generative approaches and improves both semantic understanding and computational efficiency.
@@
While most self-supervised learning objectives are designed for single-stream data, multimodal FMs that handle both vision and language require specialized approaches to manage these diverse data types. Combining multiple modalities can be handled at various stages in the FM pipeline, a spectrum known as early (or feature level) to late fusion (or decision level).
Late fusion involves using separate encoders for each data type, each with its own SSL objectives. These modalities are processed independently and combined later in the pipeline, as seen, for example, in [alayrac2022flamingo]. This method offers flexibility, allowing for the tailored processing of each modality and simplified integration at a later stage.
In contrast, early fusion integrates features from different modalities before or during the representation learning process. This results in a unified representation space that captures interactions between data types. Early fusion models, such as CLIP [radford2021learning] and ALIGN [jia2021scaling], often utilize Image-Text Contrastive Learning (ITC). ITC aligns paired images and texts by maximizing their similarity and minimizing the similarity between mismatched pairs, effectively representing related multimodal data closely in a shared embedding space. Additionally, ITC can leverage unpaired data by aligning images with separate texts describing similar concepts, using techniques such as self-distillation [bannur2023learning].
Finally, mid-fusion, and multi-layer fusion approaches, fall in between early and late fusion. These methods combine modalities at different preprocessing stages, often by creating layers within the neural network specifically designed for data fusion purposes [nagrani2021attention].
The choice between early and late fusion methods depends on the specific requirements of the task: early fusion captures fine-grained interactions between modalities, but may involve higher dimensionality. Late fusion allows for greater flexibility in unimodal encoder pipeline design, but may fail to capture complex interaction patterns. Mid fusion offers a balance by integrating modalities at specific stages, though it can lead to increased model complexity.